OpenStudy (anonymous):

Write the following using radical notation. Assume that all variables represent real positive numbers. (5b)^(4/5)

OpenStudy (anonymous):

\[(\sqrt[5]{5b})^4\]

OpenStudy (anonymous):

(5b)^(4/5) = [(5b)^4]^1/5 = (625b^4)^1/5 \[\sqrt[5]{625b ^{4}}\]

OpenStudy (anonymous):

Does that say the fifth root?

OpenStudy (anonymous):

for my answer??? yes it is the fifth root of 625b^4

OpenStudy (anonymous):

Thank you so much!